My App is a small but growing start-up that sees demand for several of its apps increasing quickly. The table below shows the last six months of downloads. Use a forecast for the first month of 215000, an initial trend forecast of 50000, and smoothing parameters of 0.05 for both demand smoothing and trend smoothing.
Month (t) Monthly Application Downloads Forecast for Next Month Trend
215,000.00 50,000.00
1 200,000
2 250,020
3 320,000
4 410,000
5 445,000
6 496,000
(Round your answer to 2 decimal places.)
a. Complete the table above, filling in the "Forecast for next month" and "Trend" columns, using double exponential smoothing.
(Round your answer to 2 decimal places.)
b. Month 7's forecast is:
Answer #1

a. Given :alpha =beta=0.05

In double exponential smoothing ,

Forecast for next month t =Forecastt-1 +Trendt-1

Given : Forecast for next month =215000 and trend of month 0 =50000

So, Forecast of month 0 =Forecast of next month-trend =215000-50000=165000

i,e In excel ,

C5=0.05*B5+0.95*(C4+D4)

(drag this cell formula for the rest of the months)

D5=0.05*(C5-C4)+0.95*D4

(Drag this cell formula for the rest of the months)

E5=C4+D4

(Drag this cell formula for the rest of the months)

Calculations are as shown in the below excel snapshot :

b. From the snapshot above,

Month 7's forecast =Forecast including trend=519273.33